# pnasresearcharticle

### v1

@@ Abstract
Antimicrobial resistance (AMR) threatens global health. A promising and underexplored strategy to tackle this problem are sequential therapies exploiting collateral sensitivity (CS), whereby resistance to one drug increases sensitivity to another. Here, we develop a four‐genotype stochastic birth–death model with two bacteriostatic antibiotics to identify switching periods that maximize bacterial extinction under subinhibitory concentrations. We show that extinction probability depends nonlinearly on switching period, with stepwise increases aligned to discrete switch events: fast sequential therapies are suboptimal as they do not allow for the evolution of resistance, a key ingredient in these therapies. A geometric distribution framework accurately predicts cumulative extinction probabilities where the per‐switch extinction probability rises with switching period. We further derive a heuristic approximation for the extinction probability based on times to fixation of single-resistant mutants. Sensitivity analyses reveal that strong reciprocal CS is required for this strategy to work, and we explore how increasing antibiotic doses and higher mutation rates modulate extinction in a nonmonotonic manner. Finally, we discuss how optimal switching periods depend on treatment duration. Our results provide quantitative design principles for in vitro and clinical sequential antibiotic therapies, underscoring the potential of CS‐guided regimens to suppress resistance evolution and eradicate infections.
keywords:
@@
Antimicrobial resistance (AMR) is rising rapidly (1), leading to higher rates of uncontrolled infections that contribute significantly to both patient mortality and healthcare costs (2). The development of new antibiotics is not fast enough to counteract this problem (3) and, although new technologies can help accelerate discovery (4), this is not guaranteed to solve the problem, as resistance to new antibiotics evolves soon after or even before their deployment in the clinic (5, 6).
@@
An alternative strategy to combat AMR is to develop multidrug treatments (7), unlocking access to large combinatorial treatment spaces. Combination therapies are the most explored alternative, where two or more antibiotics are deployed simultaneously (8). These therapies can prevent the rise of AMR (9), especially if the antibiotics are chosen ad hoc for a particular pathogen (10) or based on their interaction profiles (11). However, combination therapies are not without disadvantages: the total concentration of antibiotics introduced in the patient is higher than monotherapies, and so there is the risk of toxic effects (12). Moreover, combinations have sometimes been found to accelerate, rather than slow down, the appearance of resistance (13, 14).
@@
A less explored alternative to combination therapies is sequential therapies, in which several antibiotics are administered one after the other instead of simultaneously (15). This strategy is based on the phenomenon of collateral sensitivity (CS), in which bacterial resistance to one antibiotic increases its sensitivity to another. CS has been found in many bacterial species and antibiotic classes (16, 17, 18, 19), suggesting a promising avenue to develop treatments that can eradicate pathogenic bacterial populations (20, 21). The sequential framework opens the door to mathematical optimization approaches, where we seek the optimal sequence that maximizes the eradication of the population or minimizes the evolution of resistance (22, 23, 24, 25, 26, 27).
@@
Here, we seek antibiotic switching protocols that maximize bacterial extinction in a population model where only four genotypes and two antibiotics are considered, a common framework for both theory and experiments (28, 29). We show that sequential therapies based on strong CS can lead to bacterial eradication even at subinhibitory concentrations, provided that the correct switching period is used. The dependence of bacterial extinction on the switching period is explained by population composition, paving the way for optimization based on population metrics. Our results are contingent on the existence of strong CS, and we observe nonmonotonical relationships between extinction rates and antibiotic dose, on one hand, and mutation rates, on the other, which we can explain using our model. We end by exploring how optimal switching periods depend on treatment duration. Our work suggests that sequential therapies are a rich opportunity to explore potentially successful strategies that will help tackle the antibiotic crisis.
The model
@@
We will study bacterial population dynamics under two bacteriostatic antibiotics $A$ and $B$, using a four-genotype birth-death stochastic model. The four types are: $x_{0}$, susceptible to both antibiotics; $x_{1}$, resistant to $A$ but susceptible to $B$; $x_{2}$, resistant to $B$ but susceptible to $A$; and $x_{3}$, resistant to both (Fig. 1A). We consider mutations between types, with rates $\mu_{1}$ for the acquisition of resistance and $\mu_{2}$ for the loss of resistance. Mutations from $x_{0}$ to $x_{3}$ are not permitted, although their introduction does not qualitatively change our results. In what follows, we will refer to the population of genotype $x_{i}$ as $N_{i}$.
@@
For computational efficiency, in order to simulate the stochastic model we will use the tau-leaping algorithm (30), which approximates the dynamics of the birth-death process by taking small time increments $dt$ and generating pseudo-random Poisson-distributed numbers for all reactions: births, deaths and mutations. Population sizes are updated accordingly, and the process is repeated until desired. Note that using Gillespie’s algorithm (31), which simulates the model exactly, does not qualitatively change our results (see SI Appendix, Fig. S1, SI Appendix 1.2).
We start our simulations with antibiotic $A$ and initial population $N_{0}=k\cdot\gamma^{-1},N_{1}=N_{2}=N_{3}=0$, i.e., initially there is no resistance and the population of susceptibles is at the carrying capacity. After some elapsed time $\tau$ (switching period) we switch from antibiotic $A$ to $B$, and continue the process. At time $2\tau$ we switch back to $A$ and so on (Fig. 1B). We are interested in studying how this parameter $\tau$, the period of antibiotic switching, affects the probability that the population becomes extinct at the end of treatment.
@@
For simulation purposes, we mark a population as extinct whenever $N<0.05\gamma^{-1}$, and our results remain qualitatively unchanged by reasonable changes in this threshold (SI Appendix 1.2 and SI Appendix, Fig. S2.).
@@
Figure 2: Sequential therapies with subinhibitory antibiotic concentrations cause extinction for a wide range of switching periods. (A) Probability of extinction at the end of the treatment as switching periods vary. The intervals between two vertical lines share the same number of treatment cycles. $10,000$ trajectories were used to estimate the probability as a function of $\tau$. (B) Distribution of extinction times as a function of switching periods. Each point corresponds to the time a simulation went extinct. The colours represent the density of these events. The same background colour represents the same antibiotic used. The red line represents the end of the time we allow for switching antibiotics. (C) One thousand individual trajectories switching treatment every 20 time units. Blue, red, green, purple and orange represent the mean of trajectories that go extinct upon switching antibiotics at different swicthing events. Black represents the mean of those trajectories that do not go extinct. Dashed lines indicate the antibiotic switch. (D) Cumulative extinction probability over 10 treatment switches with different switching periods $\tau$. Points represent extinction probabilities estimated through simulation, while the solid line reflects a fit to a hierarchical geometric distribution. (E) Extiction probability for populations undergoing one antibiotic switch, blue points represent the simulation and the red points are the fitted extinction probabilities for the corresponding $\tau$ using the hierarchical geometric model (the red line is a guide to the eye). (F) Predicted extinction probabilities for populations undergoing one to four antibiotic switches using the hierarchical geometric model. Simulations were performed with final times $\tau\cdot$(numer of switches)+50+50. Parameter values are shown in SI Appendix, Table S1.
Results
@@
Sequential treatments with strong collateral sensistivity result in bacterial eradication even with subinhibitory antibiotic concentrations
@@
We start our study with subinhibitory antibiotic concentrations: $k_{A}=k_{B}=0.5$, usually called the half-maximum inhibitory concentration or IC50, and consider strong CS, $k_{\rm CS}=0.05$. This may seem like a strange place to begin, as these subinhibitory doses are usually thought to promote the evolution of resistance (32, 33). However, our simulations produce a wide range of switching periods $\tau$ that result in frequent eradication of the bacterial population. Fig. 2A shows that the probability that a population becomes extinct at the end of the treatment depends nontrivially on $\tau$. First, we observe that there are no extinctions when $\tau\to 0$ (Fig. 2A,B). As $\tau$ increases, there is a rapid increase in the probability of extinction, with two maxima at $\tau=50$ and $\tau=100$. The extinction probability shows some sharp discontinuities, which are due to a change in the number of antibiotic switches: extinction probabilities increase discontinuously when a new antibiotic switch is introduced. For instance, when $\tau$ is decreasing from $100$, the extinction probability suddenly increases when we reach $\tau=50$, where the number of antibiotic switches increases from 1 to 2. Similarly, when we cross the threshold $\tau=33.3$ there is another increase in extinction probabilities, associated with an increase in the number of switches from 2 to 3. Conversely, in the absence of switches (i.e. $\tau=0$ or $\tau>T$) there are no extinctions, which is to be expected since we are studying IC50 concentrations. Moreover, when studying extinction times, we observe that, for a fixed switching period $\tau$, extinction events occur more frequently after a short transient period following the switching time (Fig. 2B) and never before the first switch. In order to better illustrate this phenomenon, we examine a collection of trajectories for fixed $\tau=20$ (Fig. 2C): some trajectories become extinct after the first switch (marked in blue in Fig. 2C), some after the second (red), third (green), fourth (purple) and fifth (orange) switches. Populations that do not become extinct (black) eventually gain resistance.
Since extinction events occur shortly after switching antibiotics, we can visualize the extinction process as a coin-tossing game where the probability of getting heads is $p$: each time the antibiotic switches, we toss a coin. If the result is heads, the population goes extinct. The probability that the extinction event occurs exactly after $n$ switches would then be given by the geometric probability distribution $(1-p)^{n-1}p$, where $p$ is the probability that the population becomes extinct after one antibiotic switch, and which may be dependent on $\tau$. The cumulative extinction probability, i.e. the probability that the population has become extinct by the $n$th switch, is given by $1-(1-p)^{n}$. The main hypothesis of this model is that $p$ is constant throughout the process, which will be true if the time between switching is sufficiently large so that the population reaches some kind of stationary state that is independent of the switching event. Since extinction events occur shortly after switching antibiotics, we can visualize the extinction process as a coin-tossing game where the probability of getting heads is $p$: each time the antibiotic switches, we toss a coin. If the result is heads, the population goes extinct. The probability that the extinction event occurs exactly after $n$ switches would then be given by the geometric probability distribution $(1-p)^{n-1}p$, where $p$ is the probability that the population becomes extinct after one antibiotic switch, and which may be dependent on $\tau$. The cumulative extinction probability, i.e. the probability that the population has become extinct by the $n$th switch, is given by $1-(1-p)^{n}$. The main hypothesis of this model is that $p$ is constant throughout the process, which will be true if the time between switching is sufficiently large so that the population reaches some kind of stationary state that is independent of the switching event.
Fitting the cumulative extinction probability obtained from simulations for different $\tau$ to equation $1-(1-p)^{n}$ results in good agreement (SI Appendix, Fig. S3A). However, the fits are not perfect: they tend to underestimate the extinction probabilities when the number of switches is small, and vice versa. This suggests that $p$ is not constant and slightly depends on the number of switches. We capture this behavior with a hierarchical geometric model where $p=\alpha\tau+\beta n$ can change with both switching period $\tau$ and the number of switches $n$, resulting in a great fit to the data (Fig. 2D, SI Appendix 1.3). The values of $p$ obtained with our fits are shown in Fig. 2E (red line) as a function of $\tau$ and compared with the simulated extinction probabilities for treatments undergoing only one switch (blue circles): the probability of extinction per switching event is close to zero for $\tau<20$, consistent with what we observed in Fig. 2A, and increases as $\tau$ grows. In other words, the longer we wait until we switch antibiotics, the higher the probability that the population becomes extinct. The agreement between the simulated per-switch extinction (blue circles) and the hierarchical geometric fit (red lines) is quite good given the simplicity of the model. We can use this fit to predict extinction times after two or more switches, by calculating the corresponding probability using the geometric distribution, with very good agreement (Fig. 2F), supporting our hypothesis.
Heuristic explanation for the change in extinction probabilities
@@
Figure 3: Heuristic approximation for the extinction probability. (A) Sigmoid fit using the population composition before the switch. Blue circles represent the simulation; gray circles are the prediction of the sigmoid function fitted with the population before the switches; and red circles are the prediction of the sigmoid function fitted with the population before the switches for switching periods greater than 3030. (B) Heuristic fit for fixed final time treatments, considering different treatments of switching period $\tau$. (C) Heuristic fit for the extinction probability under different number of antibiotic switches. Circles are the values of extinction probabilities measured in simulations; each color represents a different number of antibiotic changes, as indicated in the legend. The solid lines indicate the estimated value of the probability of extinction for each $\tau$, using the heuristic to estimate pp. Parameter values are shown in SI Appendix, Table S1.
@@
We turn now to give an explanation for the observed extinction patterns by finding out which variables explain the dependence of the extinction probability on the switching period $\tau$. Fig. 2C hints that, in those populations that become extinct, right before the switch the population had become dominated by the single-resistant populations ($x_{1}$ or $x_{2}$) reaching a carrying capacity close to $100$. In contrast, the populations where extinction does not happen (marked in grey in Fig. 2C) are dominated by $x_{0}$, whose carrying capacity is $50$. Our hypothesis is that, for small $\tau$, resistant mutants have hardly any time to appear in the population before the antibiotic is switched and, due to the strong CS, they are rapidly invaded after the switch by type $x_{0}$, which does not go extinct under IC50 concentrations. As $\tau$ increases, however, the likelihood that either $x_{1}$ or $x_{2}$ rise in the population grows, and therefore, when the antibiotic switches, there is a chance that the whole population goes extinct.
@@
This discussion suggests that we should be able to predict extinction probabilities from the composition of the population before the switch. We have fitted the probability $p$ that a population goes extinct after an antibiotic switch to a sigmoid function $p=\big(1+e^{\mathbf{W}^{T}\mathbf{N}}\big)^{-1}$ where $\mathbf{W}$ is a parameter vector and $\mathbf{N}=(N_{0},\dots,N_{3})$ contains the population abundances right before the antibiotic switch (Fig. 3A, SI Appendix 1.4). Note that the fit is more accurate for large $\tau$ values and fails particularly when $\tau\to 0$. This was expected since more than $90\%$ of decaying populations need at least about $t=30$ time steps to become extinct before switching the antibiotic again (SI Appendix, Fig. S5B), i.e. $p$ does not only depend on population structure, and thus the sigmoid function cannot fully capture its behavior. In fact, training the sigmoid only on data where $\tau>30$ yields a more accurate fit (Fig. 3A, red circles). The fitted parameters confirm our earlier intuition (SI Appendix, Table S2): $p$ decreases when any genotype other than the single resistant increases. For example, if the population before the switch is $N_{0}=5,N_{1}=95,N_{3}=0$, then $p\approx 0.59$, but introducing one $x_{3}$ individual yields $p\approx 0.34$. This indicates that the extinction probability is largely determined by whether the population contains cells other than the currently dominant resistant genotype—the presence of even a small fraction of any other genotypes substantially decreases the chances of extinction.
@@
However, it is unrealistic to assume that complete compositional data will be available in clinical or in vitro settings to decide when to switch antibiotics. We therefore propose a heuristic approach to approximate $p$ using three key time distributions obtained from the dynamics of our four-genotype system. Specifically, we measured: (1) the distribution of times for single-resistant mutants to dominate the population (defined as exceeding $80\%$) after an antibiotic switch; (2) the distribution of extinction times following an antibiotic switch; and (3) the probability that the population contains cells other than the single-resistant mutant for the antibiotic currently in use, a choice motivated by the sigmoid fit, as previosuly discussed.
Figure 4: Extinction probability sensitivity to model parameters. (A) Collateral sensitivity (CS) is necessary for extinction. Extinction probabilities decrease as the parameter $k_{CS}$ increases, across a range of switching periods ($\tau$, shown in color). (B) Increasing antibiotic concentration (lowering $k$) robustly leads to extinction, depending on the number of treatment cycles. We observe a threshold near the MIC beyond which most populations go extinct regardless of the switching period. (C) Extinction probability as a function of antibiotic concentration for $\tau=50$. Close to the MIC, every trajectory becomes extinct. For sub-inhibitory doses, there is an intermediate dose that maximises extinction.
@@
The first two distributions were well-approximated by lognormal distributions (SI Appendix, Fig. S5), and all three can in principle be measured in vitro to obtain a practical estimate of optimal switching strategies. Building on these distributions and the previous insights gained from our geometrical model, we derived an analytical expression for the extinction probability $p$ as a function of $\tau$:
@@
where $p_{d}(\tau)$ is the probability that a decaying population goes extinct within time $\tau$ (SI Appendix, Fig. S5B), and $p_{r}(\tau)$ is the probability that a single-resistant mutant is not completely dominant in the population, given by
@@
with $p_{\mathrm{x_{1}}}(\tau)$ denoting the probability that genotype $x_{1}$ dominates the population at time $\tau$, given that the system started with $x_{2}$ in dominance (SI Appendix, Fig. S5A). Intuitively, for short times the initial condition $N_{0}(0)=50$ provides a good approximation, while for longer times the condition $N_{2}(0)=90$ becomes more accurate. The weighting between these two scenarios is naturally captured by the distribution of takeover times of the resistant strain, represented by $p_{\mathrm{x_{1}}}(\tau)$.
@@
We use the geometric distribution formula to extend these extinction probabilities to various antibiotic changes. However, some care has to be taken with the initial and boundary conditions at the end of the treatment (SI Appendix 1.5). This formula captures our previous intuitions for the two necessary conditions for extinction: the dominance of the single-resistant genotype and the existence of long enough decay times. This simple heuristic matches the shape of extinction probabilities derived from full stochastic simulations (Fig. 3B) and gives an accurate estimate for cumulative extinction probabilities of populations under one, two or more antibiotic switches (Fig. 3C).
@@
Additional scenarios: weak collateral sensitivity, increasing antibiotic concentrations and changing mutation rates

### v3

# Optimization of sequential therapies to maximize extinction of resistant bacteria through collateral sensitivity

@@ Abstract
Antimicrobial resistance (AMR) threatens global health. A promising and underexplored strategy to tackle this problem are sequential therapies exploiting collateral sensitivity (CS), whereby resistance to one drug increases sensitivity to another. Here, we develop a four‐genotype stochastic birth–death model with two bacteriostatic antibiotics to identify switching periods that maximize bacterial extinction under subinhibitory concentrations. We show that extinction probability depends nonlinearly on switching period, with stepwise increases aligned to discrete switch events: fast sequential therapies are suboptimal as they do not allow for the evolution of resistance, a key ingredient in these therapies. A geometric distribution framework accurately predicts cumulative extinction probabilities, where the per‐switch extinction probability rises with switching period. We further derive a heuristic approximation for the extinction probability based on times to fixation of single-resistant mutants. Sensitivity analyses reveal that strong reciprocal CS is required for this strategy to work, and we explore how increasing antibiotic doses and higher mutation rates modulate extinction in a nonmonotonic manner. Finally, we discuss how longer therapies maximize extinction but also cause higher resistance, leading to a Pareto front of optimal switching periods. Our results provide quantitative design principles for in vitro and clinical sequential antibiotic therapies, underscoring the potential of CS‐guided regimens to suppress resistance evolution and eradicate infections.
@@ I Introduction
Antimicrobial resistance (AMR) is rising rapidly [1], leading to higher rates of uncontrolled infections that contribute significantly to both patient mortality and healthcare costs [2]. The development of new antibiotics is not fast enough to counteract this problem [3] and, although new technologies can help accelerate discovery [4], this is not guaranteed to solve the problem, as resistance to new antibiotics evolves soon after or even before their deployment in the clinic [5, 6].
@@
An alternative strategy to combat AMR is to develop multidrug treatments [7], unlocking access to large combinatorial treatment spaces. Combination therapies are the most explored alternative, where two or more antibiotics are deployed simultaneously [8]. These therapies can prevent the rise of AMR [9], especially if the antibiotics are chosen ad hoc for a particular pathogen [10] or based on their interaction profiles [11]. However, combination therapies are not without disadvantages: the total concentration of antibiotics introduced in the patient is higher than monotherapies, and so there is the risk of toxic effects [12]. Moreover, combinations have sometimes been found to accelerate, rather than slow down, the appearance of resistance [13, 14].
@@
A less explored alternative to combination therapies is sequential therapies, in which several antibiotics are administered one after the other instead of simultaneously [15]. This strategy is based on the phenomenon of collateral sensitivity (CS), in which bacterial resistance to one antibiotic increases its sensitivity to another. CS has been found in many bacterial species and antibiotic classes [16, 17, 18, 19], suggesting a promising avenue to develop treatments that can eradicate pathogenic bacterial populations [20, 21]. The sequential framework opens the door to mathematical optimization approaches, where we seek the optimal sequence that maximizes the eradication of the population or minimizes the evolution of resistance [22, 23, 24, 25, 26, 27].
@@
Here, we seek antibiotic switching protocols that maximize bacterial extinction in a population model where only four genotypes and two antibiotics are considered, a common framework for both theory and experiments [28, 29]. We show that sequential therapies based on strong CS can lead to bacterial eradication even at subinhibitory concentrations, provided that the correct switching period is used. The dependence of bacterial extinction on the switching period is explained by population composition, paving the way for optimization based on population metrics. Our results are contingent on the existence of strong CS, and we observe nonmonotonic relationships between extinction rates and antibiotic dose, on one hand, and mutation rates, on the other, which we can explain using our model. We end by searching for optimal switching periods that both maximize extinction and minimize the appearance of resistance, finding there is a trade-off between both objectives, leading to a Pareto front of optimal therapies. Our work suggests that sequential therapies are a rich opportunity to explore potentially successful strategies that will help tackle the antibiotic crisis.
II The model
@@
We will study bacterial population dynamics under two bacteriostatic antibiotics $A$ and $B$, using a four-genotype birth-death stochastic model. The four types are: $x_{0}$, susceptible to both antibiotics; $x_{1}$, resistant to $A$ but susceptible to $B$; $x_{2}$, resistant to $B$ but susceptible to $A$; and $x_{3}$, resistant to both (Fig. 1(a)). We consider stochastic transitions (mutations) between types, with rates $\mu_{1}$ for the acquisition of resistance and $\mu_{2}$ for the loss of resistance [30]. Mutations from $x_{0}$ to $x_{3}$ are not permitted, although their introduction does not qualitatively change our results. In what follows, we will refer to the population of genotype $x_{i}$ as $N_{i}$.
@@
For computational efficiency, in order to simulate the stochastic model we will use the tau-leaping algorithm [31], which approximates the dynamics of the birth-death process by taking small time increments $dt$ and generating pseudo-random Poisson-distributed numbers for all reactions: births, deaths and mutations. Population sizes are updated accordingly, and the process is repeated until desired. Note that using Gillespie’s algorithm [32], which simulates the model exactly, does not qualitatively change our results (see Appendix A, Fig. 8).
We start our simulations with antibiotic $A$ and initial population $N_{0}=k\cdot\gamma^{-1},N_{1}=N_{2}=N_{3}=0$, i.e., initially there is no resistance and the population of susceptible is at the carrying capacity. After some elapsed time $\tau$ (switching period) we switch from antibiotic $A$ to $B$, and continue the process. At time $2\tau$ we switch back to $A$ and so on (Fig. 1(b)). We are interested in studying how this parameter $\tau$, the period of antibiotic switching, affects the probability that the population becomes extinct at the end of treatment.
@@
For simulation purposes, we mark a population as extinct whenever $N<0.05\gamma^{-1}$ [33], and our results remain qualitatively unchanged by reasonable changes in this threshold (Appendix A and Appendix A, Fig. 9).
@@
Figure 2: Sequential therapies with subinhibitory antibiotic concentrations cause extinction for a wide range of switching periods. (a) Probability of extinction at the end of the treatment as switching periods vary. The intervals between two vertical lines share the same number of treatment cycles. $10,000$ trajectories were used to estimate the probability as a function of $\tau$. (b) Distribution of extinction times as a function of switching periods. Each point corresponds to the time a simulation went extinct. The colors represent the density of these events. The same background color represents the same antibiotic used. The red line represents the end of the time we allow for switching antibiotics. (c) One thousand individual trajectories switching treatment every 20 time units. Blue, red, green, purple and orange represent the mean of trajectories that go extinct upon switching antibiotics at different switching events. Black represents the mean of those trajectories that do not go extinct. Dashed lines indicate the antibiotic switch. (d) Cumulative extinction probability over 10 treatment switches with different switching periods $\tau$. Points represent extinction probabilities estimated through simulation, while the solid line reflects a fit to a hierarchical geometric distribution. (e) Extinction probability for populations undergoing one antibiotic switch, blue points represent the simulation and the red points are the fitted extinction probabilities for the corresponding $\tau$ using the hierarchical geometric model (the red line is a guide to the eye). (f) Predicted extinction probabilities for populations undergoing one to four antibiotic switches using the hierarchical geometric model. Simulations were performed with final times $\tau\cdot$(number of switches)+50+50. Parameter values in Appendix B, Table 1.
III Results
@@
III.1 Sequential treatments with strong collateral sensitivity result in bacterial eradication even with subinhibitory antibiotic concentrations
@@
We start our study with subinhibitory antibiotic concentrations: $k_{A}=k_{B}=0.5$, usually called the half-maximum inhibitory concentration or IC50, and consider strong CS, $k_{\rm CS}=0.05$. This may seem like a strange place to begin, as these subinhibitory doses are usually thought to promote the evolution of resistance [34, 35]. However, our simulations produce a wide range of switching periods $\tau$ that result in frequent eradication of the bacterial population. Fig. 2(a) shows that the probability that a population becomes extinct at the end of the treatment depends non-trivially on $\tau$. First, we observe that there are no extinctions when $\tau\to 0$ (Fig. 2(a,b)). As $\tau$ increases, there is a rapid increase in the probability of extinction, with two maxima at $\tau=50$ and $\tau=100$. The extinction probability shows some sharp discontinuities, which are due to a change in the number of antibiotic switches: extinction probabilities increase discontinuously when a new antibiotic switch is introduced. For instance, when $\tau$ is decreasing from $100$, the extinction probability suddenly increases when we reach $\tau=50$, where the number of antibiotic switches increases from 1 to 2. Similarly, when we cross the threshold $\tau=33.3$ there is another increase in extinction probabilities, associated with an increase in the number of switches from 2 to 3. Conversely, in the absence of switches (i.e. $\tau=0$ or $\tau>T$) there are no extinctions, which is to be expected since we are studying IC50 concentrations. Moreover, when studying extinction times, we observe that, for a fixed switching period $\tau$, extinction events occur more frequently after a short transient period following the switching time (Fig. 2(b)) and never before the first switch. In order to better illustrate this phenomenon, we examine a collection of trajectories for fixed $\tau=20$ (Fig. 2(c)): some trajectories become extinct after the first switch (marked in blue in Fig. 2(c)), some after the second (red), third (green), fourth (purple) and fifth (orange) switches. Populations that do not become extinct (black) eventually gain resistance.
@@
Figure 3: Heuristic approximation for the extinction probability. (a) Sigmoid fit for the extinction probability at the end of the treatment, using the population composition before the switch. Blue circles represent the simulation; gray circles are the prediction of the sigmoid function fitted with the population before the switches; and red circles are the prediction of the sigmoid function fitted with the population before the switches for switching periods greater than $30$. (b) Heuristic fit for the same probability as in (a) using fixed final time treatments, and considering several switching periods $\tau$. (c) Heuristic fit for the extinction probability under different number of antibiotic switches. Circles are the values of extinction probabilities measured in simulations; each color represents a different number of antibiotic changes, as indicated in the legend. The solid lines indicate the estimated value of the probability of extinction for each $\tau$, using the heuristic to estimate pp. Parameter values in Appendix B, Table 1.
Since extinction events occur shortly after switching antibiotics, we can visualize the extinction process as a coin-tossing game where the probability of getting heads is $p$: each time the antibiotic switches, we toss a coin. If the result is heads, the population goes extinct. The probability that the extinction event occurs exactly after $n$ switches would then be given by the geometric probability distribution $(1-p)^{n-1}p$, where $p$ is the probability that the population becomes extinct after one antibiotic switch, and which may be dependent on $\tau$. The cumulative extinction probability, i.e. the probability that the population has become extinct by the $n$th switch, is given by $1-(1-p)^{n}$. The main hypothesis of this model is that $p$ is constant throughout the process, which will be true if the time between switching is sufficiently large so that the population reaches some kind of stationary state that is independent of the switching event.
Fitting the cumulative extinction probability obtained from simulations for different $\tau$ to equation $1-(1-p)^{n}$ results in good agreement (Appendix C, Fig. 10(a)). However, the fits are not perfect: they tend to underestimate the extinction probabilities when the number of switches is small, and vice versa. This suggests that $p$ is not constant and slightly depends on the number of switches. We capture this behavior with a hierarchical geometric model where $p=\alpha\tau+\beta n$ can change with both switching period $\tau$ and the number of switches $n$, resulting in a great fit to the data (Fig. 2(d), Appendix C). The values of $p$ obtained with our fits are shown in Fig. 2(e) (red line) as a function of $\tau$ and compared with the simulated extinction probabilities for treatments undergoing only one switch (blue circles): the probability of extinction per switching event is close to zero for $\tau<20$, consistent with what we observed in Fig. 2(a), and increases as $\tau$ grows. In other words, the longer we wait until we switch antibiotics, the higher the probability that the population becomes extinct. The agreement between the simulated per-switch extinction (blue circles) and the hierarchical geometric fit (red lines) is quite good, given the simplicity of the model. We can use this fit to predict extinction times after two or more switches, by calculating the corresponding probability using the geometric distribution, with very good agreement (Fig. 2(f)), supporting our hypothesis.
III.2 Heuristic explanation for the change in extinction probabilities
@@
We turn now to give an explanation for the observed extinction patterns by finding out which variables explain the dependence of the extinction probability on the switching period $\tau$. Fig. 2(c) hints that, in those populations that become extinct, right before the switch the population had become dominated by the single-resistant populations ($x_{1}$ or $x_{2}$) reaching a carrying capacity close to $100$. In contrast, the populations where extinction does not happen (marked in gray in Fig. 2(c)) are dominated by $x_{0}$, whose carrying capacity is $50$. Our hypothesis is that, for small $\tau$, resistant mutants have hardly any time to appear in the population before the antibiotic is switched and, due to the strong CS, they are rapidly invaded after the switch by type $x_{0}$, which does not go extinct under IC50 concentrations. As $\tau$ increases, however, the likelihood that either $x_{1}$ or $x_{2}$ rise in the population grows, and therefore, when the antibiotic switches, there is a chance that the whole population goes extinct.
Figure 4: Distributions for key transition times in the model. Simulated distributions (blue histograms) and fitted lognormal probability density functions (dashed curves) for two temporal processes underlying the heuristic extinction estimate. (a) Time until the system transitions between stable states following an antibiotic change. Starting from $x_{2}(0)=80$ and $x_{i}(0)=0$ for $i=0,1,3$ under antibiotic AA until $x_{1}(t)>80$. (b) Time to extinction under a new antibiotic, measured only for simulations where extinction occurs. (c) Time dependent probability of having $x_{0}+x_{2}+x_{3}\geq 2$ starting from the initial conditions indicated by colors in the legend. $10000$ trajectories were simulated for estimating these distributions.
@@
This discussion suggests that we should be able to predict extinction probabilities from the composition of the population before the switch. We have fitted the probability $p$ that a population goes extinct after an antibiotic switch to a sigmoid function $p=\big(1+e^{\mathbf{W}^{T}\mathbf{N}}\big)^{-1}$ where $\mathbf{W}$ is a parameter vector and $\mathbf{N}=(N_{0},\dots,N_{3})$ contains the population abundances right before the antibiotic switch (Fig. 3(a), Appendix D). Note that the fit is more accurate for large $\tau$ values and fails particularly when $\tau\to 0$. This was expected since more than $90\%$ of decaying populations need at least about $t=30$ time steps to become extinct before switching the antibiotic again (Fig. 4(b)), i.e. $p$ does not only depend on population structure, and thus the sigmoid function cannot fully capture its behavior. In fact, training the sigmoid only on data where $\tau>30$ yields a more accurate fit (Fig. 3(a), red circles). The fitted parameters confirm our earlier intuition (Appendix D, Table 2): $p$ decreases when any genotype other than the single resistant increases. For example, if the population before the switch is $N_{0}=5,N_{1}=95,N_{3}=0$, then $p\approx 0.59$, but introducing one $x_{3}$ individual yields $p\approx 0.34$. This indicates that the extinction probability is largely determined by whether the population contains cells other than the currently dominant resistant genotype—the presence of even a small fraction of any other genotypes substantially decreases the chances of extinction.
@@
However, it is unrealistic to assume that complete compositional data will be available in clinical or in vitro settings to decide when to switch antibiotics. We therefore propose a heuristic approach to approximate $p$ using three key time distributions obtained from the dynamics of our four-genotype system. Specifically, we measured: (1) the distribution of times for single-resistant mutants to dominate the population (defined as exceeding $80\%$) after an antibiotic switch; (2) the distribution of extinction times following an antibiotic switch; and (3) the probability that the population contains cells other than the single-resistant mutant for the antibiotic currently in use, a choice motivated by the sigmoid fit, as previously discussed.
@@
The first two distributions were well-approximated by lognormal distributions (Fig. 4), and all three can in principle be measured in vitro to obtain a practical estimate of optimal switching strategies. Building on these distributions and the previous insights gained from our geometric model, we derived an analytical expression for the extinction probability $p$ as a function of $\tau$:
@@
where $p_{d}(\tau)$ is the probability that a decaying population goes extinct within time $\tau$ (Fig. 4(b)), and $p_{r}(\tau)$ is the probability that a single-resistant mutant is not completely dominant in the population, given by
@@
with $p_{\mathrm{x_{1}}}(\tau)$ denoting the probability that genotype $x_{1}$ dominates the population at time $\tau$, given that the system started with $x_{2}$ in dominance (Fig. 4(a)). Intuitively, for short times the initial condition $N_{0}(0)=50$ provides a good approximation, while for longer times the condition $N_{2}(0)=90$ becomes more accurate. The weighting between these two scenarios is naturally captured by the distribution of takeover times of the single-resistant strain, represented by $p_{\mathrm{x_{1}}}(\tau)$.
Figure 5: Extinction probability sensitivity to model parameters. (a) Collateral sensitivity (CS) is necessary for extinction. Extinction probabilities decrease as the parameter $k_{CS}$ increases, across a range of switching periods ($\tau$, shown in color). (b) Increasing antibiotic concentration (lowering $k$) robustly leads to extinction, depending on the number of treatment cycles. We observe a threshold near the MIC, beyond which most populations go extinct regardless of the switching period. (c) Extinction probability as a function of antibiotic concentration for $\tau=50$. Close to the MIC, every trajectory becomes extinct. For subinhibitory doses, there is an intermediate dose that maximizes extinction.
@@
We use the geometric distribution formula to extend these extinction probabilities to various antibiotic changes. However, some care has to be taken with the initial and boundary conditions at the end of the treatment (Appendix E). This formula captures our previous intuitions for the two necessary conditions for extinction: the dominance of the single-resistant genotype and the existence of long enough decay times. This simple heuristic matches the shape of extinction probabilities derived from full stochastic simulations (Fig. 3(b)) and gives an accurate estimate for cumulative extinction probabilities of populations under one, two or more antibiotic switches (Fig. 3(c)).
@@
III.3 Additional scenarios: weak collateral sensitivity, increasing antibiotic concentrations and changing mutation rates